# high-definition

### v1

@@ Abstract
Until open-world foundation models match the performance of specialized approaches, the effectiveness of deep learning models remains heavily dependent on dataset availability. Training data must align not only with the target object categories but also with the sensor characteristics and modalities. To bridge the gap between available datasets and deployment domains, domain adaptation strategies are widely used. In this work, we propose a novel approach to transfer sensor-specific knowledge from an image dataset to LiDAR, an entirely different sensing domain. Our method leverages detections on camera images to create a semantic parametric map. The map elements are modeled to produce pseudo labels in the target domain without any manual annotation effort. Unlike previous domain transfer approaches, our method does not require direct overlap between sensors and enables extending the angular perception range from a front-view camera to a full $360\text{\,}\mathrm{\SIUnitSymbolDegree}$ view. On our large-scale road feature dataset, XD-MAP outperforms single shot baseline approaches by +19.5 mIoU for 2D semantic segmentation, +19.5 $\textrm{PQ}_{\textrm{th}}$ for 2D panoptic segmentation, and +32.3 mIoU in 3D semantic segmentation. The results demonstrate the effectiveness of our approach achieving strong performance on LiDAR data without any manual labeling.
@@ 2 Related Work
Due to their relevance for automated driving, road signs have been detected and mapped early on [30, 37], already using planes in 3D space as representation. Traffic lights are commonly detected in camera images since this allows estimating their lights’ state [8, 17]. Usually, they are mapped as point with an optional height or a number of lights. Purely geometric primitives, e.g. edges [40], flat surfaces [40, 15], and poles [29, 15], have been proposed as compact yet robust semantic features for SLAM and localization and are often extracted from stereo cameras or lidar point clouds. They are represented as planes or lines, occasionally with a diameter for poles.
@@ 4.1 XD-MAP
Our 3D point cloud annotation utilizes the same intermediate annotation and combines each 2D shape with an object-specific depth, creating a frustum that encapsulates the corresponding map element. All LiDAR points falling within this frustum are assigned with the label of the map element. This method ensures that the 3D labels accurately reflect the semantic information from the map while maintaining spatial consistency within the point cloud.
@@ 4.2 Baseline Approaches
We implement two single shot baselines, XD-B1 and XD-B2. The first approach specializes in preserving the object’s shape, while the second aims to retain the precise LiDAR depth for each point within each object. The key challenge these approaches address is how to accurately infer both the 3D position and spatial extent of an object in LiDAR space when it is initially represented as a 2D polygon in the camera image. Given that the camera image provides a limited field of view, the baselines can only provide annotations for roughly $100\text{\,}\mathrm{\SIUnitSymbolDegree}$. Within this region, object instances detected in the camera image are projected into LiDAR space and annotated accordingly, while non-object points are labeled as background to clearly distinguish them. Additionally, points or pixels outside the supervised FOV are annotated to be ignored during training to avoid negatively affecting the optimization process.
@@ 5.1 Dataset, Task Definition and Metrics
Task Definition and Metric
We evaluate our cross-modal domain adaptation system with the three tasks of 2D semantic segmentation, 3D semantic segmentation and panoptic segmentation. This covers common computer vision task modalities and the common input forms of Lidar data, range images and 3D point clouds. The evaluation regimes follow established definitions and metrics. For 2D semantic segmentation we report the IoU as calculated in [7] and for 3D semantic segmentation we follow the task definition and evaluation metric of Behley et al. [4]. Panoptic segmentation is evaluated with the panoptic segmentation quality as defined in [13]. All metrics are calculated with the three object classes of poles, traffic lights and traffic signs, dropping the background class which is present during training. For panoptic segmentation this results in reporting only the metrics for the ”thing” classes, as the background is the only ”stuff” class in the dataset.
@@ 5.2 Models and Hyperparameter Setting
For the task of 3D segmentation, we chose the Cylinder3D [42] architecture. We adopt the hyperparameters for training on SemanticKITTI [34], except for the number of epochs. The model is trained 12 epochs with batch size of 16. AdamW optimizer is used with a base learning rate of 0.001 and weight decay of 0.01.
@@ 5.4 Results
XD-MAP shows strong performance increases of +27.2 and +19.5 mIoU points in the 2D semantic segmentation task compared to the single shot baselines XD-B1 and XD-B2, respectively. Similar increases are consistent across all element classes. Metrics for panoptic segmentation mirror this increase, with an improvement of +21.7 and +19.5 $\textrm{PQ}_{\textrm{th}}$ to XD-B1 and XD-B2. Looking at the other metrics, it can be seen that this gain can almost entirely be attributed to the gain in recognition quality $\textrm{PQ}_{\textrm{th}}$, meaning models trained with XD-MAP have much higher precision and recall of detected elements. These results clearly demonstrate the benefits of a systematic approach of a highly accurate SLAM in combination with parametric mapping of geometric primitives for cross-modal domain adaptation in contrast with simpler approaches limited by camera field of view or inaccuracies of the pseudo-labels.
@@ 6 Limitations
While we have demonstrated XD-MAP for domain adaptation from camera to lidar, we argue that it is feasible for many sensor combinations, environments, and semantic classes. This comes with a few limitations: While we do not need an overlapping field of view of the sensors, it is required that both sensors perceive the same objects over time. Also the resolution of both sensors need to be adequate to the size, level of detail and maximum distance of mapped objects. Additionally, as we aim for self-supervised domain adaptation, by the nature of the task we have no manually annotated ground truth available. Since the baselines only provide a small field of view and visual inspection shows the much better accuracy of XD-MAP annotations, these were selected as our ground truth. The XD-MAP pipeline also requires a highly accurate SLAM and multimodal calibration. Sensor suites with very noisy sensors or without accurate calibration may lack the prerequisites for semantic parametric mapping of instances. Furthermore, objects that are not representable with geometric primitives are not able to be processed with the proposed system, which may limit its applicability outside of structured domains like roads in autonomous driving.
@@ 7 Conclusion
We prove the effectiveness of our approach by transferring knowledge from panoptically labeled RGB images from a front-view camera to a $360\text{\,}\mathrm{\SIUnitSymbolDegree}$ lidar as target. Evaluating three complementary tasks, XD-MAP outperforms the best baseline by +19.5 mIoU for 2D semantic segmentation, +19.5 $\textrm{PQ}_{\textrm{th}}$ for 2D panoptic segmentation, and +32.3 mIoU in 3D semantic segmentation. This shows that the method generalizes well across a comprehensive array of typical tasks. For applications, XD-MAP hence makes the knowledge, which is captured in the broad variety of semantically rich and well-generalizing image datasets, such as Mapillary Vistas, available to still underrepresented or specialized sensors, such as lidar. While we demonstrated XD-MAP with three exemplary classes of static objects, we cannot wait to extend XD-MAP to more semantic classes, including dynamic objects that require 4D reconstruction, as well as more sensors, such as imaging radar.

### v2

@@ Abstract
Until open-world foundation models match the performance of specialized approaches, deep learning systems remain dependent on task- and sensor-specific data availability. To bridge the gap between available datasets and deployment domains, domain adaptation strategies are widely used. In this work, we propose XD-MAP, a novel approach to transfer sensor-specific knowledge from an image dataset to LiDAR, an entirely different sensing domain. Our method leverages detections on camera images to create a semantic parametric map. The map elements are modeled to produce pseudo labels in the target domain without any manual annotation effort. Unlike previous domain transfer approaches, our method does not require direct overlap between sensors and enables extending the angular perception range from a front-view camera to a full $360\text{\,}\mathrm{\SIUnitSymbolDegree}$ view. On our large-scale road feature dataset, XD-MAP outperforms single shot baseline approaches by +19.5 mIoU for 2D semantic segmentation, +19.5 $\textrm{PQ}_{\textrm{th}}$ for 2D panoptic segmentation, and +32.3 mIoU in 3D semantic segmentation. The results demonstrate the effectiveness of our approach achieving strong performance on LiDAR data without any manual labeling.
@@ 2 Related Work
Due to their relevance for automated driving, road signs have been detected and mapped early on [30, 37], already using planes in 3D space as representation. Traffic lights are commonly detected in camera images since this allows estimating their lights’ state [8, 17]. Usually, they are mapped as a point with an optional height or a number of lights. Purely geometric primitives, e.g. edges [40], flat surfaces [40, 15], and poles [29, 15], have been proposed as compact yet robust semantic features for SLAM and localization and are often extracted from stereo cameras or LiDAR point clouds. They are represented as planes or lines, occasionally with a diameter for poles.
@@ 4.1 XD-MAP
Our 3D point cloud annotation utilizes the same intermediate annotation and combines each 2D shape with an object-specific depth, creating a frustum that encapsulates the corresponding map element. All LiDAR points falling within this frustum are assigned the label of the map element. This method ensures that the 3D labels accurately reflect the semantic information from the map while maintaining spatial consistency within the point cloud.
@@ 4.2 Baseline Approaches
We implement two single-shot baselines, XD-B1 and XD-B2. The first approach specializes in preserving the object’s shape, while the second aims to retain the precise LiDAR depth for each point within each object. The key challenge these approaches address is how to accurately infer both the 3D position and spatial extent of an object in LiDAR space when it is initially represented as a 2D polygon in the camera image. Given that the camera image provides a limited field of view, the baselines can only provide annotations for roughly $100\text{\,}\mathrm{\SIUnitSymbolDegree}$. Within this region, object instances detected in the camera image are projected into LiDAR space and annotated accordingly, while non-object points are labeled as background to clearly distinguish them. Additionally, points or pixels outside the supervised FOV are marked to be ignored during training to avoid negatively affecting the optimization process.
@@ 5.1 Dataset, Task Definition and Metrics
Task Definition and Metrics
We evaluate our cross-modal domain adaptation system with the three tasks of 2D semantic segmentation, 3D semantic segmentation and panoptic segmentation. This covers common computer vision task modalities and the common input representations of LiDAR data, range images and 3D point clouds. The evaluation regimes follow established definitions and metrics. For 2D semantic segmentation we report the IoU as calculated in [7] and for 3D semantic segmentation we follow the task definition and evaluation metric of Behley et al. [4]. Panoptic segmentation is evaluated with the panoptic segmentation quality as defined in [13]. All metrics are calculated with the three object classes of poles, traffic lights and traffic signs, dropping the background class which is present during training. For panoptic segmentation this results in reporting only the metrics for the ”thing” classes, as the background is the only ”stuff” class in the dataset.
@@ 5.2 Models and Hyperparameter Setting
For the task of 3D segmentation, we chose the Cylinder3D [42] architecture. We adopt the hyperparameters for training on SemanticKITTI [34], except for the number of epochs. The model is trained for 12 epochs with a batch size of 16. AdamW optimizer is used with a base learning rate of 0.001 and weight decay of 0.01.
@@ 5.4 Results
XD-MAP shows strong performance increases of +27.2 and +19.5 mIoU points in the 2D semantic segmentation task compared to the single shot baselines XD-B1 and XD-B2, respectively. Similar increases are consistent across all element classes. Metrics for panoptic segmentation mirror this increase, with an improvement of +21.7 and +19.5 $\textrm{PQ}_{\textrm{th}}$ compared to XD-B1 and XD-B2. Looking at the other metrics, it can be seen that this gain can almost entirely be attributed to the gain in recognition quality $\textrm{PQ}_{\textrm{th}}$, meaning models trained with XD-MAP have much higher precision and recall of detected elements. These results clearly demonstrate the benefits of a systematic approach of a highly accurate SLAM in combination with parametric mapping of geometric primitives for cross-modal domain adaptation in contrast with simpler approaches limited by camera field of view or inaccuracies of the pseudo-labels.
@@ 6 Limitations
While we have demonstrated XD-MAP for domain adaptation from camera to LiDAR, we argue that it is applicable for many sensor combinations, environments, and semantic classes. This comes with a few limitations: While we do not need an overlapping field of view of the sensors, it is required that both sensors perceive the same objects over time. Also the resolution of both sensors needs to be adequate to the size, level of detail and maximum distance of mapped objects. Since the baselines only provide a small field of view and visual inspection shows the much better accuracy of XD-MAP annotations, these were selected as our ground truth. The XD-MAP pipeline also requires a highly accurate SLAM and multimodal calibration. Sensor suites with very noisy sensors or without accurate calibration may lack the prerequisites for semantic parametric mapping of instances. Furthermore, objects that are not representable with geometric primitives are not able to be processed with the proposed system, which may limit its applicability outside of structured domains like roads in autonomous driving.
@@ 7 Conclusion
We prove the effectiveness of our approach by transferring knowledge from panoptically labeled RGB images from a front-view camera to a $360\text{\,}\mathrm{\SIUnitSymbolDegree}$ LiDAR as target. Evaluating three complementary tasks, XD-MAP outperforms the best baseline by +19.5 mIoU for 2D semantic segmentation, +19.5 $\textrm{PQ}_{\textrm{th}}$ for 2D panoptic segmentation, and +32.3 mIoU in 3D semantic segmentation. This shows that the method generalizes well across a comprehensive array of typical tasks. For applications, XD-MAP hence makes the knowledge, which is captured in the broad variety of semantically rich and well-generalizing image datasets, such as Mapillary Vistas, available to still underrepresented or specialized sensors, such as LiDAR. While we demonstrated XD-MAP with three exemplary classes of static objects, future work will extend XD-MAP to more semantic classes, including dynamic objects that require 4D reconstruction, as well as more sensors, such as imaging radar.
Acknowledgment
The research leading to these results was partially funded by the German Federal Ministry for Economic Affairs and Energy (BMWE) and the European Union within the just better DATA (jbDATA) project, grant number 19A23003H. The authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg. HoreKa is partly funded by the German Research Foundation (DFG).